A chemist reacts 45.0 mL of 3.4 M AgNO3 with an excess of Na2CrO4. How many grams of silver chromate will be produced?


1
Expert's answer
2022-03-09T07:13:05-0500

Solution:

Molarity = Moles of solute / Liters of solution

Therefore,

Moles of AgNO3 = Molarity of AgNO3 × Liters of AgNO3 solution

Moles of AgNO3 = (3.4 M) × (45.0 mL) × (1 L / 1000 mL) = 0.153 mol


Balanced chemical equation:

2AgNO3 + Na2CrO4 → Ag2CrO4 + 2NaNO3

According to stoichiometry:

2 mol of AgNO3 produces 1 mol of Ag2CrO4

Thus, 0.153 mol of AgNO3 produces:

Moles of Ag2CrO4 = (0.153 mol AgNO3) × (1 mol Ag2CrO4 / 2 mol AgNO3) = 0.0765 mol Ag2CrO4


The molar mass of Ag2CrO4 is 331.73 g/mol

Therefore,

Mass of Ag2CrO4 = (0.0765 mol Ag2CrO4) × (331.73 g Ag2CrO4 / 1 mol Ag2CrO4) = 25.377 g Ag2CrO4

Mass of Ag2CrO4 = 25.377 g Ag2CrO4 = 25.4 g Ag2CrO4


Answer: 25.4 grams of silver chromate (Ag2CrO4) will be produced